22. A body weighing 1kg is thrown vertically upwards at a speed of 10m / s. The potential energy at the highest point will be: A) 1000 B) 500 C) 250 D) OJ


1
Expert's answer
2021-05-31T08:45:19-0400

When the body reaches the highest point its kinetic energy converted into potential energy:


"PE=KE=\dfrac{1}{2}mv^2,""PE=KE=\dfrac{1}{2}\cdot1\ kg\cdot(10\ \dfrac{m}{s})^2=50\ J."

None of the answers are correct. The correct answer is 50 J.